OpenStudy (anonymous):

The normal 120V rms 60Hz line voltage is half-wave-rectified and connected to a clipper circuit that clips the wave so that it will never exceed 120V. What is the average value of the voltage?

OpenStudy (anonymous):

peak voltage Vp=Vrmsx1.414 Vp=120x1.414= 169.68 average value or DC= Vp/Pi DC=169.68x 3.1416= 54.01 volt dc

OpenStudy (anonymous):

oops sorry DC= 169.68/3.1416= 54.01 volts